# SMARTe-VR: Student Monitoring and Adaptive Response Technology for e-learning in Virtual Reality

### v1

@@ 2.3 Public State of the Art Database
Public datasets in education mainly focus on student responses and assessment (see Table 1), with limited incorporation of biometric data and without VR. For example, the ASSISTments (ASSISTments Foundation 2013) and Junyi (Chang et al. 2015) datasets provide responses to math questions. Similarly, EdNet dataset (Choi et al. 2020) includes question responses and incorporates records of materials consulted by students during TOEIC exam preparation. All of these datasets lack both biometric and VR monitoring.
@@
Few datasets explore multimodality with biometrics. The SAD-IRT dataset (Zhou, Suzuki, and Kumano 2024) includes facial tracking during literacy tasks, while edBB (Hernandez-Ortega et al. 2020a) records biometric signals such as heart rate, EEG and facial video for behavior analysis. However, these datasets focus on PC-based assessment.
@@
In contrast, SMARTe-VR uniquely combines VR, learning, and assessment, with biometric monitoring, making it a valuable resource for AL system development.
3 SMARTe-VR: Platform for Adaptive E-Learning in Virtual Reality
SMARTe-VR platform is designed for e-learning sessions through VR, enabling instructors to upload their educational content while students experience it in an immersive environment that simulates a lecture hall with a screen, similar to works such as (Nagao 2023). The platform provides a VR environment that facilitates student interaction with the system, monitors learning data, and captures facial expressions in real-time. This feature allows detailed analysis of the student’s emotional and cognitive state. Additionally, it integrates learning assessment tools, enabling instructors to adjust content based on student performance.
3.1 Platform Features
VR Interaction: The immersive environment simulates a virtual lecture hall, allowing students to zoom in/out by moving their head, adjust brightness, and interact using VR controllers or a mouse. Features include answering ques- tions, opening textbook, highlighting text, and controlling video playback (pause, fast-forward, etc.).
Function Color
Marks learned content. Yellow
Highlights content for review. Blue
Indicates content not understood. Red
Marks key content for memorization. Green
Table 2: Functions for tagging and highlighting with their associated Colors in SMARTe-VR
Structure and Assessment Tool: Instructors can structure LSs using lecture videos of various lengths and content. It also features an automatic assessment system (Auto QA) that can be activated at different stages. Before the session: Preliminary questions assess initial student knowledge. After each video: Comprehension of the video content is assessed.
This method, used by platforms like Coursera and edX, is supported by research showing that a few well-crafted questions effectively measure real-time understanding (Cummins, Beresford, and Rice 2015). These assessments support adaptive models such as IRT and KT, enabling the platform to be used for intelligent AL.
Student Feedback: After each Auto QA intervention, students receive their results and a comparison with the top-performing student. At the session’s end, a score summary is provided. Continuous feedback and performance comparisons motivate students, helping them track progress, identify focus areas, and optimize learning using notes and resources (Van der Kleij, Feskens, and Eggen 2015).
  • Textbook Highlighting: Students can mark text with a color-coded system (see Table 2) indicating importance and comprehension level.
  • Quick Tagging Buttons: Four colour-coded buttons (see Table 2) enable students to tag video sections quickly. The platform labels these sections, providing valuable insights for instructors and learning analysis.
Studies show that visual interaction mechanisms improve content retention and comprehension. Highlighting key information is associated with higher retention and more efficient review (Ogata et al. 2022).
@@ 3.1 Platform Features
Table 3: Overview of devices and data collected by the monitoring system.
3.2 SMARTe-VR Database [^1] Access: https://github.com/BLANCELIN1/SMARTe-VR-DB
We monitored LSs focused on the TOEIC, specifically Section 5, which assesses grammar comprehension. The TOEIC is a recognized English test used by companies to evaluate non-native speakers’ language skills in professional contexts. This study involved 10 Japanese students from Hiroshima University. The dataset comprises 450 answered questions and approximately 25 hours of learning data, including facial movements and expressions.
Ethical Considerations.
The database was collected in accordance with the Declaration of Helsinki and approved by the Ethics Committee. Learners were briefed on the procedure and data acquisition, providing informed consent. Due to the extended monitoring, participants received an incentive, with an additional reward for top performers.
Devices.
@@
VR Headset: We used the VIVE XR Elite, designed for educational and business applications, along with the VIVE XR Facial Tracker. This accessory uses infrared cameras to measure interpupillary distance for optimal viewing and enable facial tracking. Using the OpenXR SDK, we monitored up to 51 facial features at 30 Hz, ensuring system stability and preventing overheating. Captured facial features include movements and expressions, quantified from 0 to 1 in intensity. These features include eye movements and blinking, jaw opening, mouth and lip expressions (e.g., smiles) cheek gestures (e.g., puffing), and tongue movements.
@@
Personal Computer: It connected to the VIVE XR Elite manages the virtual environment, collects and analyzes sensor data, and captures learning records, ensuring optimal performance without headset overheating. Although VIVE XR Elite controllers are supported, we chose the mouse as the VR controller because: 1) Students remained seated, simulating a classroom, making it more practical; 2) Test feedback indicated it was more comfortable for interacting with the SMARTe-VR interface and highlighting text; 3) This approach allows for better comparisons to traditional e-learning setups for future research.
Protocol and Tasks
@@
The goal was to design a monitoring system for real VR LSs, covering both learning and assessment processes while tracking learning metadata and facial expressions. This protocol aims to aid the research community in developing frameworks like IRT, KT, MMLA, and BBLR, filling the gap in databases with these features.
@@
The content of each lecture video and its questions was designed to create a balanced dataset, trying an equal number of videos were labeled as understood or not understood. Ten 8-minute TOEIC exam lecture videos, explained in Japanese, were selected from YouTube and varied in difficulty according to TOEIC guidelines. Each video had three multiple-choice questions with four options, focused on sentence completion to assess grammar and vocabulary. A video was labeled understood if a student answered at least 2 out of 3 questions correctly. The questions were from a TOEIC dataset on the ProProfs platform, which includes data on response accuracy and participant numbers.
@@
Using the sample size estimation formula, we determined the minimum number of responses needed to represent 99% of the population with a 1% margin of error, enabling the classification of questions by difficulty and content. Questions were categorized as easy (80% correct response rate), medium (50%), and difficult (20%). The question structure per video was: easy-level videos had 2 easy and 1 medium question; medium-level had 3 medium questions; difficult-level had 2 difficult and 1 medium question.
To verify difficulty accuracy, a preliminary test with similar students was conducted, collecting subjective evaluations on each session’s difficulty and its questions.
@@
Before Starting the Session:
@@
Calibration and Student Information: The supervisor explained the session protocol and verified data capture. A 3D facial model was monitored throughout to confirm recording of facial movements and expressions. Privacy Policy: The students reviewed a document detailing the data collection and its purpose, with the option to accept or decline participation. Video Lecture 0 and Auto QA Trial Phase: The students began with an initial lecture video that explained how to use the interface. At the end of this video, the Auto QA system was activated, presenting 5 questions without a time limit to help students become familiar with the interface. Auto QA Pretest: A pretest of 10 questions (3 easy, 4 medium, and 3 difficult) related to the upcoming lecture content was designed with two main objectives: i) to evaluate the students’ current knowledge level and ii) to serve as a baseline for IRT or similar systems.
Learning Session:
@@
The session consisted of 10 lecture videos, each followed by 3 related questions. The structure was: i) Students watched an 8-minute video, interacting with the textbook and color-coded buttons; ii) The Auto QA was activated post-video, presenting 3 questions with a one-minute response time each, allowing students to review the textbook. This immediate questioning reduced potential memory issues; iii) Students received feedback and compared their performance to the top student; iv) Students could take a break if needed—most did not, but those who did paused for approximately 5 minutes; v) A 5-10 minute VR headset battery change occurred after the 5th video.
End of the Session:
@@
The students received a total score and a comparison with the top-performing student. They provided feedback to the supervisor about their SMARTe-VR experience, which highlighted the following: i) Prolonged VR use did not cause motion sickness as participants remained seated in the VR; ii) Students reported restlessness over time due to limited interaction options; and iii) Some students reported discomfort when using the mouse in VR, as they were unable to see their hands.We conclude that integrating Mixed Reality (MR) and shortening sessions could enhance interaction, device usability (e.g., mouse, keyboard), and overall user experience.
Challenges
Figure 1: Example setup and data captured by the SMARTe-VR platform, including the proposed understanding detection models (SMART-TCN and SMART-MLP) for database validation.
@@
We propose 5 challenges for the research community, each with specific target and input data. The objective is to leverage the SMARTe-VR database to develop AL models that estimate student performance, knowledge, or behavior. The proposed challenges include:
@@
Challenge 1 - Understanding Estimation: Predicting student comprehension of educational content is valuable for AL. We propose estimating if a student understood lecture content using binary labels (understood, not understood). Target: Binary classification to predict student understanding in VR based on multimodal data. Inputs: Facial features, metadata, pretest results, and content difficulty.
@@
Challenge 3 - Student Feedback Prediction: The ability to anticipate how a student perceives content (e.g., important, learned/not learned, to memorize) is crucial for AL. This challenge seeks to create a model that uses color-coded button interactions from VR sessions as ground truth to predict feedback. Target: Predict student feedback on their understanding and the perceived relevance of the content. Inputs: Facial features, learning metadata, pretest results, and lecture videos.
@@
Challenge 4 - Response Time Detection: Predicting if a student will respond before the time limit can help gauge confidence, detect fatigue or inattention, and adjust workload or response time dynamically. Only biometric data and metadata from the first 20 seconds are permitted. Target: Predict if a student will respond on time. Inputs: Facial features, metadata (e.g., selection path history), content and question difficulty, and pretest results.
@@
Challenge 5 - Performance Prediction: Student performance was measured by accuracy (percentage of correct responses), enabling classification into three levels (low, medium, high). We propose estimating final performance across three levels using biometric data and session metadata without incorporating response data. Indicators such as attention, emotional state, and fatigue from facial features, along with learning metadata, can help in predicting performance. Target: Performance. Inputs: Facial features, learning metadata, content difficulty, and pretest results.
4 Experiments and Results
4.1 IRT Models
@@
We tested two IRT (1PL) model architectures based on (Zhou, Suzuki, and Kumano 2024) to validate the dataset, aiming to predict student understanding during video lectures (Challenge 1). These models used facial features within a time window ${W_{l}\in\{1,\ldots,8\}}$ minutes, along with one-hot vectors for the user and the questions. The binary understand label served as the ground truth for prediction. Fig. 1 shows a schematic of these models:
@@
SMART-TCN model: Based on Deep-IRT (see Section 3), we used two twin Multilayer Perceptrons (MLPs) to estimate question difficulty ($b$) and student ability ($\theta$) separately. One MLP processed a one-hot vector representing the question to obtain $b$, while the other processed a one-hot vector identifying the student to estimate $\theta$. The student state ($\phi$) was based on SAD-IRT, using a TCN that processed facial features extracted from video lectures via the VR headset. The TCN was configured with 4 layers (32 to 64 neurons), a filter size of 3, and an expansion rate of 1, 1, 2, and 4, optimized to capture subtle changes in expressions held for extended periods. The TCN output was then connected to a fully connected layer with an output layer using a Tanh function to obtain the state parameter. Finally, the system estimated the probability of understanding as $P(\text{Understand})=\frac{1}{1+e^{-(\theta-(b+\phi))}}$.
@@
SMART-MLP Model: This model was similar to the SMART-TCN but replaced the TCN with three parallel MLPs to estimate each parameter independently. Each MLP had two hidden layers, the first with 128 neurons and the second with 64 neurons, both using ReLU activation, followed by an output layer with a Tanh function. This model used the global facial features as input to estimate $\phi$.
@@
For SMART-TCN model, we used local facial features, with the VR headset providing 30 sets of 51 features per second. The resulting vector is $\textbf{f}_{\mathrm{L}}\in\mathbb{R}^{30\times 51\times W_{l}}$, dependent on the selected time window $W_{l}$.
Global Features
@@
For SMART-MLP model, we used global facial features, specifically 8 global features for each facial feature. These are the maximum, minimum, mean, standard deviation, median, kurtosis, skewness, and rate of change. The resulting vector is $\textbf{f}_{\mathrm{G}}\in\mathbb{R}^{8\times 51}$.
4.2 Experimental Protocol
SMARTe-VR is a nearly balanced dataset regarding conference understand labels, with 55% of sessions marked as understood. Specifically, 83% of easy, 45% of medium, and 40% of difficult sessions are marked as understood. Each user has 10 labels for each learning session, totaling 100 labels across all users.
We used a “leave-one-out” cross-validation approach, where each user is left out once for testing while the remaining users are used for training; this process repeats for each user. However, even with leave-one-out, the model is always trained on video lecture 0 and the 10 pretest questions, including those from the left-out user, to estimate their initial ability values. The decision threshold is set at the point where the False Positive and False Negative rates are equal.
@@ 4.2 Experimental Protocol
Figure 2: Comparison of accuracy across different time windows ($W_{l}$) for our proposed models and the Traditional IRT model on the SMARTe-VR database.
4.3 Results

### v3

@@ 2.3. Existing Research Datasets
Public data sets in education focus mainly on student responses and assessment (see Table 1), with limited incorporation of biometric data and without VR. For example, the ASSISTments (ASSISTments Foundation, 2013) and Junyi (Chang et al., 2015) data sets provide responses to math questions. Similarly, the EdNet dataset (Choi et al., 2020) includes question responses and incorporates records of materials consulted by students during the preparation for the TOEIC exam. All of these data sets lack biometric and VR monitoring.
@@
Few datasets explore multimodality with biometrics. The SAD-IRT dataset (Zhou et al., 2024) includes facial tracking during literacy tasks, while edBB (Hernandez-Ortega et al., 2020a) records biometric signals such as heart rate, EEG, and facial video for behavior analysis. Similarly, the IMPROVE (Daza et al., 2025a) dataset captures multimodal biometric and behavioral data during short e-learning sessions, with assessment being the main focus of data collection. IMPROVE was specifically designed to investigate the effects of mobile phone usage on student learning (Becerra et al., 2025b). However, these data sets focus on PC-based assessment.
@@
In contrast, SMARTe-VR uniquely combines VR, learning, and assessment with biometric monitoring, making it a valuable resource for the development of AL systems.
3. SMARTe-VR: Platform Description
Table 2. Overview of devices and data collected by the monitoring system.
@@ 3. SMARTe-VR: Platform Description
SMARTe-VR platform is designed for VR-based e-learning sessions, allowing instructors to upload their educational content while students experience it in an immersive environment that simulates a lecture hall with a screen, similar to works such as (Nagao, 2023). The platform provides a virtual reality environment that facilitates student interaction with the system, monitors learning data, and captures facial expressions in real time. This feature allows for a detailed analysis of the student’s emotional and cognitive state. In addition, it integrates learning assessment tools, allowing instructors to adjust content based on student performance.
3.1. Platform Features
VR Interaction: The immersive environment simulates a virtual lecture hall, allowing students to zoom in/out by moving their head, adjust brightness, and interact using VR controllers or a mouse. Features include answering questions, opening the textbook, highlighting text, and controlling video playback (pause, fast-forward,  etc.).
Structure and Assessment Tool: Instructors can structure LSs using lecture videos of various lengths and content. It also features an automatic assessment system (Auto QA) that can be activated at different stages. Before the session: Preliminary questions assess initial student knowledge. After each video: Comprehension of the content is assessed. This method, used by platforms such as Coursera and edX, is supported by research showing that a few well-crafted questions effectively measure real-time understanding (Cummins et al., 2015). These assessments support models such as IRT and KT, enabling the platform to be used for intelligent AL.
Student Feedback: After each Auto QA intervention, students receive their results and a comparison with the highest performing student. At the end of the session, a summary of the scores is provided. Continuous feedback and performance comparisons motivate students, helping them track progress, identify focus areas, and optimize learning using notes and resources (Van der Kleij et al., 2015).
  • Textbook Highlighting: Students can mark text using a four-color coding system, where each color represents a distinct function: learned content, material for review, lack of understanding, and key content for memorization.
  • Quick Tagging Buttons: A set of four color-coded buttons allows students to quickly tag video sections, using the same meaning as in text highlighting. The platform labels these sections, providing valuable information for instructors and learning analytics.
Studies show that visual interaction mechanisms improve content retention and comprehension. Highlighting key information is associated with higher retention and more efficient review (Ogata et al., 2022).
3.2. Contributed Database: SMARTe-VR-db
We monitor LSs focused on the TOEIC, specifically Section 5, which assesses grammar comprehension. The TOEIC is a recognized English test used by companies to evaluate nonnative speakers’ language skills in professional contexts. This study involved 10 Japanese students from Hiroshima University. The data set comprises 450 answered questions and approximately 25 hours of learning data, including facial movements and expressions. The SMARTe-VR-db is publicly available on GitHub (Daza et al., 2025b).
Ethical Considerations: The database was collected in accordance with the Declaration of Helsinki and approved by the Ethics Committee. The learners were briefed about the procedure and data acquisition, providing informed consent. Due to extended monitoring, participants received an incentive with an additional reward for top performers.
3.2.1. Devices
@@
VR Headset: We used the VIVE XR Elite, designed for educational and business applications, together with the VIVE XR Facial Tracker. This accessory uses infrared cameras to measure interpupillary distance for optimal viewing and facial tracking. Using the OpenXR SDK, we monitored up to 51 facial features at 30 Hz, ensuring system stability and preventing overheating. Captured facial features include movements and expressions, quantified from 0 to 1 intensity. These features include eye movements and blinking, jaw opening, mouth and lip expressions (e.g. smiles), cheek gestures (e.g. puffing) and tongue movements. The VIVE XR Facial Tracker combines dual infrared cameras for eye movement tracking (120 Hz) and a monochrome camera (60 Hz) with an ultra-wide 151° field of view to capture facial movements below the eyes, enabling reliable detection of subtle expressions even under low-light conditions. The combination of a wide-angle field of view, inertial sensors, and facial cameras embedded in the VR headset ensures stable tracking during head motion, full rotations, or suboptimal device alignment (typical limitations of standard webcams used in traditional e-learning setups). The immersive VR context further promotes natural and expressive facial behavior, improving the fidelity of facial biometrics (Dubovi, 2024).
@@
Personal Computer: Connected to the VIVE XR Elite, it managed the virtual environment, collected and analyzed sensor data, and recorded learning logs while maintaining optimal performance without overheating the headset. Although VIVE XR Elite controllers are supported, we chose the mouse as the VR controller because: 1) students remained seated, simulating a classroom, making it more practical; 2) test feedback indicated it was more comfortable to interact with the SMARTe-VR interface and highlight text; 3) this approach allows for better comparisons to traditional e-learning setups for future research.
3.2.2. Protocol and Tasks
@@
The goal was to design a monitoring system for real VR LSs, covering both learning and assessment processes while tracking learning metadata and facial expressions. This protocol aims to help the research community in developing frameworks like IRT, KT, MMLA, and BBLR, filling the gap in research data sets with these features.
@@
The content of each lecture video and its questions was designed to create a balanced dataset, with an equal number of videos labeled understood or not understood. Ten 8-minute TOEIC exam lecture videos, explained in Japanese, were selected from YouTube and varied in difficulty according to the TOEIC guidelines. Each video had three multiple-choice questions with four options, focused on sentence completion to assess grammar and vocabulary. A video was labeled understood if a student answered at least 2 out of 3 questions correctly. The questions were from a TOEIC data set on the ProProfs platform, which includes data on the precision of the response and the number of participants.
@@
Using sample size estimation, we determined the minimum number of responses needed to represent 99% of the population with a 1% margin of error, allowing the classification of questions by difficulty and content. The questions were classified as easy (80% correct response rate), medium (50%), and difficult (20%). The question structure per video was the following: easy-level videos had 2 easy and 1 medium question; medium-level had 3 medium questions; difficult-level had 2 difficult and 1 medium question. To verify difficulty, a preliminary test was conducted with similar students, collecting subjective evaluations on the difficulty of each session and its questions.
@@
Figure 2. Example setup and data captured by the SMARTe-VR platform, including the proposed understanding detection models (SMART-TCN and SMART-MLP) for database validation. Both follow a Deep IRT-based architecture (Tsutsumi et al., 2021) for estimating student ability and item difficulty from one-hot encodings of users and questions. They also use facial features to estimate the student state, which is their only point of difference: SMART-TCN applies a TCN to local facial features, while SMART-MLP uses an MLP on global facial features.
[…]
The figure shows the setup of the SMARTe-VR platform, the data captured during the learning sessions, and how this information is used by two proposed IRT-based models—SMART-TCN and SMART-MLP—to predict student understanding during video lectures.
@@
Before Starting the Session: Calibration and Student Information: The supervisor explained the protocol and verified data capture. A 3D facial model was monitored to confirm the recording of facial movements and expressions. Privacy Policy: The students reviewed a document detailing data collection and its purpose, with the option to accept or decline participation. Video Lecture 0 and Auto QA Trial Phase: The students began with an initial lecture video explaining how to use the interface. At the end of this video, the Auto QA system was activated, presenting 5 questions without a time limit to help students become familiar with the interface. Auto QA Pretest: A pre-test of 10 questions (3 easy, 4 medium, and 3 difficult) related to the upcoming lecture content was designed with two main objectives: i) to evaluate the current level of knowledge of the students and ii) to serve as a baseline for IRT or similar systems.
@@
Learning Session: The session consisted of 10 lecture videos, each followed by 3 related questions. The structure was as follows: i) Students watched an 8-minute video, interacting with the textbook and color-coded buttons. ii) Auto-QA was activated after the video, presenting 3 questions with a one-minute response time each, allowing students to review the textbook. This immediate questioning reduced potential memory problems. iii) Students received feedback and compared their performance to the top student. iv) Students could take a break if needed; most did not, but those who did paused for approximately 5 minutes. And v) a 5–10 minute battery change of the VR headset occurred after the fifth video.
@@
End of the Session: The students received a total score and a comparison with the highest performing student. They provided feedback to the supervisor about their SMARTe-VR experience, which highlighted the following: i) prolonged VR use did not cause motion sickness as participants remained seated in the VR, ii) students reported restlessness over time due to limited interaction options, and iii) some students reported discomfort when using the mouse in VR, as they were unable to see their hands. We conclude that integrating Mixed Reality (MR) and shortening sessions could enhance interaction, device usability (e.g. mouse, keyboard) and overall user experience.
3.2.3. Challenges
@@
We propose 5 challenges for the research community, each with a specific target and input data. The objective is to take advantage of the SMARTe-VR database to develop AL models that estimate student performance, knowledge, or behavior. The proposed challenges are as follows.
@@
Challenge 1 - Understanding Estimation: Predicting student comprehension of educational content is valuable for AL. We propose estimating whether a student understood the lecture content using binary labels (understood, not understood). Target: Binary classification to predict student understanding in VR based on multimodal data. Inputs: Facial features, metadata (e.g. highlighting, button presses), pre-test results and content difficulty.
@@
Challenge 3 - Student Feedback Prediction: The ability to anticipate how a student perceives content (e.g., important, learned/not learned, to memorize) is crucial for AL models. This challenge seeks to create a model that uses color-coded button interactions from VR sessions as ground truth to predict feedback. Target: Predict student feedback on their understanding and perceived relevance of the content. Inputs: Facial features, learning metadata, pretest results, and lecture videos.
@@
Challenge 4 - Response Time Detection: Predicting if a student will respond before the time limit can help gauge confidence, detect fatigue or inattention, and dynamically adjust workload or response time. Only biometric data and metadata from the first 20 seconds are allowed. Target: Predict if a student will respond on time. Inputs:  Facial features, metadata (e.g., selection path history), content and question difficulty, and pretest results.
@@
Challenge 5 - Performance Prediction: Student performance was measured by accuracy (percentage of correct responses), allowing classification into three levels (low, medium, high). We propose estimating the final performance across three levels using biometric data and session metadata without incorporating response data. Indicators such as attention, emotional state, and fatigue of facial features, along with learning metadata, can help predict performance. Target: Performance. Inputs: Facial features, learning metadata, content difficulty, and pretest results.
4. Experiments and Results
4.1. IRT Models
@@
We tested two IRT (1PL) models based on (Zhou et al., 2024) to validate our methods for Challenge 1: to predict student understanding during video lectures. These models use facial features in a time window $W_{l}\in\{1,\ldots,8\}$} (in minutes), along with one-hot vectors for the user and the questions. The binary understand label served as the ground truth for prediction. Fig. 2 shows a schematic of these  models:
@@
SMART-TCN model: Based on Deep-IRT (see Section 2), we used two twin Multilayer Perceptrons (MLPs) to estimate the difficulty of the question ($b$) and the ability of the student ($\theta$) separately. An MLP processed a one-hot vector representing the question to obtain $b$, while the other processed a one-hot vector identifying the student to estimate $\theta$. The student state ($\phi$) was based on SAD-IRT, using a TCN that processed facial features extracted from video lectures using the VR headset. The TCN was configured with 4 layers (32 to 64 neurons), a filter size of 3, and an expansion rate of 1, 1, 2, and 4, optimized to capture subtle changes in expressions held for extended periods. The TCN output was then connected to a fully connected layer with an output layer using a Tanh function to obtain the state parameter. Finally, the system estimated the probability of understanding as:
(1) $P(\text{Understand})=\frac{1}{1+e^{-(\theta-(b+\phi))}}$
@@
SMART-MLP Model: This model was similar to SMART-TCN but replaced the TCN with an MLP for student state estimation, resulting in a structure with three parallel MLPs to estimate each parameter independently. Each MLP had 2 hidden layers, the first with 128 neurons and the second with 64 neurons, both using ReLU activation, followed by an output layer with a Tanh function. This model used global facial features as input to estimate $\phi$.
@@
Local Features. For the SMART-TCN model, we used local facial features, with the VR headset providing 30 sets of 51 features per second. The resulting vector is $\textbf{f}_{\mathrm{L}}\in\mathbb{R}^{30\times 51\times W_{l}\times 60}$, with $W_{l}$ being the time window in minutes.
@@
Global Features. For the SMART-MLP model, we used 8 global statistics for each facial feature. These are the maximum, minimum, mean, standard deviation, median, kurtosis, skewness, and rate of change. The resulting vector is $\textbf{f}_{\mathrm{G}}\in\mathbb{R}^{8\times 51}$.
Traditional IRT baseline. To compare the proposed models, we implemented a traditional IRT model (1PL, Rasch) as a baseline, using the following estimation process. First, item difficulties were approximated using a sign-inverted Probit transformation of the empirical accuracy rates as initial values. Then, the abilities of the students were estimated by MLE by maximizing the likelihood of their observed responses. Both item difficulties and student abilities were updated in an iterative process until convergence (Baker, 2001).
@@
Figure 3. Comparison of accuracy across different time windows ($W_{l}$) for our proposed models and the Traditional IRT baseline on the SMARTe-VR database.
[…]
Bar chart showing the accuracy obtained on the SMARTe-VR database for three models: SMART-TCN, SMART-MLP, and the Traditional IRT baseline.
4.2. Experimental Protocol
SMARTe-VR is a nearly balanced data set on understanding labels, with 55% of sessions marked as understood. Specifically, 83% of easy, 45% of medium, and 40% of difficult sessions are marked as understood. Each user has 10 labels for each learning session, totaling 100 labels for all users.
We used a leave-one-out cross-validation approach, where each user is left out once for testing while the remaining users are used for training; this process repeats for each user. However, even with leave-one-out, the model is always trained on video lecture 0 and the 10 pretest questions, including those from the left-out user, to estimate their initial ability values. The decision threshold is set at the point where the False Positive and False Negative rates are equal.
4.3. Results